# DenseSplat: Densifying Gaussian Splatting SLAM with Neural Radiance Prior

### v1

@@ Abstract
Index Terms:
Visual Dense SLAM, Neural Radiance Field, 3D Gaussian Splatting
@@ I Introduction
Visual Dense Simultaneous Localization and Mapping (SLAM) is a core area of study in 3D computer vision, focusing on real-time localization of the camera and generating a fine-grain map of the surrounding environment. It plays a crucial role in robot localization and navigation, autonomous vehicles, and Virtual/Augmented Reality (VR/AR).
Recent breakthroughs in differential rendering, specifically with Neural Radiance Fields (NeRF) [1, 2] and 3D Gaussian Splatting (3DGS) [3, 4, 5], have significantly advanced visual dense SLAM systems [6, 7, 8, 9, 10, 11]. NeRF-based neural implicit SLAM [12, 13, 14, 15, 16] integrates the NeRF model for simultaneous tracking and mapping, which facilitates high-quality, dense online map reconstruction and remarkably improves geometric accuracy. Building on this, Gaussian-based SLAM systems [17, 18, 19, 20, 21, 22, 23] have pushed the boundaries by delivering higher-fidelity map reconstruction. By utilizing explicit Gaussian primitives, 3DGS provides substantial benefits in terms of detailed texture representation [18], explicit scene manipulation [24], and remarkable real-time rendering capabilities [23].
@@
However, 3DGS faces particular challenges within the context of SLAM systems. Unlike offline reconstruction often applied to exhaustive image collections [13, 25], online SLAM systems are typically deployed in complex environments with insufficient observations, leading to partially observed or obstructed views. This shortage significantly affects the completeness of the Gaussian map, as it struggles to effectively interpolate missing geometry of unobserved areas. Moreover, current Gaussian SLAM systems rely on per-frame pixel sampling from the input stream, which fails to capture the 3D structure of the scene. Such 2D-based initialization results in a map representation where foreground elements are detailed with denser primitives, while the backgrounds, encompassing broader scene geometries, are less defined. This disparity creates a challenging optimization landscape for multiview geometry [26]. Given these limitations, current Gaussian SLAM systems often maintain a dense keyframe list, such as one in every four frames, during the mapping process that attempts to engage more viewpoints for reliable map optimization [27]. This reliance on a dense keyframe list demands extensive memory and reduces real-time processing capabilities, both of which are crucial for online SLAM systems.
To address these challenges, we propose employing NeRF priors to densify the Gaussian SLAM system. This densification is achieved through two primary mechanisms using the NeRF model: (1) its interpolation capabilities efficiently close the gaps in the map with densely positioned Gaussian primitives, and (2) it offers a robust initialization of Gaussian primitives that can be densified using extremely sparse keyframes. Moreover, NeRF-based sampling ensures an even distribution of Gaussian primitives aligned with the scene geometry and allows for manageable granularity in the scene representation through specific sampling ratios. Coupled with loop closure and bundle adjustment (BA), our method delivers state-of-the-art performance across large-scale synthetic and real-world datasets. Overall, our contributions can be summarized as follows:
@@ II-A Neural Dense SLAM
In contrast to traditional SLAM systems [28, 29, 30, 31, 32, 33, 34] that utilize point-clouds or voxels for sparse map representation, neural dense SLAM systems [35, 36, 37, 38, 39, 40, 41, 42, 43] offer substantial advantages through their dense neural radiance maps, providing a robust foundation for downstream tasks in robotics and AR/XR [44]. iMAP [45] pioneered neural implicit SLAM but suffers from large tracking and mapping error using a single MLP. NICE-SLAM [46] employs multiple MLPs for coarser-to-finer mapping, effectively filling the gaps in reconstruction. ESLAM [47] leverages tri-plane features for efficient scene representation, while Co-SLAM [48] employs multi-resolution hash-grids for real-time performance. Point-SLAM [49] relies on neural point clouds for dense scene reconstruction, and Loopy-SLAM [50] introduces map corrections to address scene drift caused by accumulated tracking errors. However, maps reconstructed using NeRF typically lack the quality seen in more recent systems built upon 3D Gaussian Splatting. Furthermore, high-resolution NeRF models often require extensive training and exhibit slower real-time rendering performance [3], which significantly reduces their practical efficiency.
@@ II-B Gaussian Splatting SLAM
Propelled by advancements in 3DGS [3], recent Gaussian SLAM systems [17, 27, 24, 23, 51] have shown remarkable capabilities in high-fidelity map reconstruction and efficient real-time rendering. Notably, SplaTAM [18] utilizes isotropic Gaussian representation coupled with dense point-cloud sampling to ensure geometric precision. Conversely, MonoGS [19] employs anisotropic Gaussians to accelerate map reconstruction and enhance texture rendering. Nevertheless, the frame-to-frame tracking mechanisms of these systems do not incorporate loop closure or bundle adjustment, which leads to significant tracking discrepancies in real-world environments. Concurrently, Photo-SLAM [20] and RTG-SLAM [21], which integrate feature-based tracking [29, 31] with dense Gaussian maps, achieve superior tracking accuracy and real-time performance. However, they take the trade-off of diminished rendering quality due to the sparse sampling of scene representation. Moreover, systems such as Gaussian-SLAM [52] and LoopSplat [53] propose to implement submap division and fusion strategies to tackle the high memory consumption of fine-grained Gaussian maps. While Gaussian SLAM systems offer superior rendering quality compared to NeRF models, the explicit and discrete nature of their scene representation often results in substantial gaps and holes in the reconstructed map due to unobserved or obstructed views commonly seen in online systems. These deficiencies severely affect their efficiency in real-world applications.
@@ II-C Gaussian Splatting with Neural Radiance Prior
Vanilla 3DGS [3] utilizes the sparse point cloud derived from Structure-from-Motion [54, 55], whose reconstruction quality heavily relies on the accuracy of the initial point cloud [56]. As an alternative, RadSplat [26] introduces NeRF model [1] into the Gaussian Splatting framework, aiming to achieve robust real-time rendering of complex real-world scenes. RadSplat [26] specifically employs NeRF as a prior for initialization and supervision, enabling fast training convergence and enhanced quality.
Compared to RadSplat [26], our method differs fundamentally in task objectives and motivations. RadSplat [26] focuses on offline reconstruction scenarios where datasets typically consist of 360-degree or dense viewpoint coverage [57, 25], making completeness and rendering quality the primary goals. In contrast, our method addresses the challenges of real-time SLAM systems deployed in robotic systems, where the limited sensor field of view and navigation path lead to sparse and sequential observations. These factors frequently result in unobserved or partially observed geometry due to obstacles, posing significant challenges for explicit Gaussian representations that often leave critical gaps in the reconstruction. To overcome these issues, our method utilizes NeRF not solely for initialization but as a robust mechanism for interpolating unobserved regions and enabling real-time adaptability in environments with sparse views. These challenges are typically absent in offline scenarios.
@@ III-A Neural Radiance Prior
Preliminaries of Neural Radiance Rendering The NeRF model $f_{\theta}$ [1] is a continuous function that predicts colors $\mathbf{c}\in\mathbb{R}^{3}$ and volume density $\sigma\in\mathbb{R}_{+}$ along the sampled rays $\mathbf{r}$. Specifically, given a camera origin $\mathbf{o}\in\mathbb{R}^{3}$ and ray direction $\mathbf{v}$, by uniformly sampling $\mathrm{M}$ points $\mathbf{x}=\mathbf{o}+\mathbf{d}_{i}\mathbf{v}|_{i\in\{1...\mathrm{M}\}}$, the pixel color $\mathbf{C}_{\mathrm{N}}$ can be rendered using ray marching:
$\mathbf{C}_{\mathrm{N}}=\sum_{i=1}^{M}\mathbf{c}_{i}\alpha_{i}T_{Ni},\quad\text{and}\quad T_{Ni}=\prod_{k=1}^{i-1}(1-\alpha_{k})~{},$ (1)
@@
NeRF-based Camera Tracking We track the camera-to-world transformation matrix $\mathbf{T}_{wc}\in\mathrm{SE}(3)$ for each frame $t$ by optimizing the pose using the objective function of NeRF defined in Equation 2. The pose is initialized based on the principle of constant velocity:
$\mathbf{T}_{t}=\mathbf{T}_{t-1}\mathbf{T}^{-1}_{t-2}\mathbf{T}_{t-1}~{}.$ (3)
Here, we employ the NeRF model exclusively for per-frame camera tracking, after which we apply loop closure and bundle adjustment based on the fine-grained Gaussian map to correct accumulative drift, detailed in Section III-C.
Geometry-aware Point Sampling Compared to explicit Gaussian representations that require dense viewpoints for thorough scene optimization, NeRF-based models offer remarkable advantages in terms of interpolation, as they can deduce unseen geometry [58, 46, 59]. To harness this capability for efficient real-time performance, we utilize a multi-resolution hash radiance field as proposed by [60]. This grid system enables precise interrogation of volumetric data across various resolutions, effectively capturing detailed surface geometries even in sparsely sampled areas. We identify critical surface transitions by setting a density threshold, $\tau_{\rm grid}$, which examines each grid edge to detect transitions where the density at one corner surpasses $\tau_{\rm grid}$, while at its adjacent corner, it does not. At these transitions, we interpolate the surface crossing points using:
@@ III-B Fine-grained Gaussian Map
Multi-scale Gaussian Rendering By initializing through grid sampling from th NeRF model, we represent the scene using a set of anisotropic 3D Gaussian primitives ${\{\mathcal{G}_{j}\}}$ [3]. Each primitive $\mathcal{G}_{j}$ is defined by a mean $\mu_{j}\in\mathbb{R}^{3}$, a covariance matrix $\Sigma_{j}$, an opacity value $\alpha_{j}\in[0,1]$, coefficients for third-order spherical harmonics $\mathbf{SH}_{j}\in\mathbb{R}^{16}$, and a scaling factor $s_{j}\in\mathbb{R}^{3}$. During the rendering process, these primitives are first projected onto a 2D plane, transforming them into 2D Gaussians. The transformation utilizes a viewing matrix $W$, and the resulting 2D covariance matrix $\Sigma^{\prime}_{j}$ in the image space is computed as follows [61]:
@@
The radiance field enables interpolated point initialization for Gaussian primitives, yet it can also lead to aliasing effects that diminish the quality of the map during sampling [13, 62]. This issue becomes particularly acute at the edges of objects where small Gaussian floaters result in significant artifacts. Drawing inspiration from [63], we adopt a multi-scale Gaussian rendering strategy that consolidates smaller Gaussians into larger ones to enhance scene consistency. Specifically, we use Gaussian functions across four levels of detail, corresponding to down-sampling resolutions of 1×, 4×, 16×, and 64×. Throughout the training phase, we merge smaller, fine-level Gaussians into larger, coarse-level Gaussians. The selection of Gaussians for merging is determined by pixel coverage, which picks Gaussians based on the coverage range set by the inverse of the highest frequency component in the region, denoted as $f_{\rm max}=1/s_{j}$, where $s_{j}$ is the scaling factor.
Ray-Guided Gaussian Pruning Sampling from the NeRF model may also introduce erroneous Gaussian floaters and outliers that affect the quality of the reconstruction. To reduce redundant Gaussians produced by the densification process and enhance rendering efficiency, we implement a pruning strategy based on sampled rays from the NeRF model. Specifically, we use an importance assessment to identify and remove inactive Gaussians from the map during optimization. The importance of each Gaussian is quantified based on its contribution to sampled rays across all input images $\{I_{i}\}_{i=1}^{N}$. Drawing inspiration from [26], we implement the score function for each primitive as:
$\mathcal{E}(\mathcal{G}_{i})=\max_{r\in I_{f}}(\alpha_{i}^{r}T_{\mathcal{G}i}^{r})~{},$ (7)
where $\alpha_{i}^{r}T_{Gi}^{r}$ captures the Gaussian $\mathcal{G}_{i}$’s contribution to the final color prediction of a pixel along the ray $r$. We then compute a pruning mask:

### v2

@@ I Introduction
Visual Dense Simultaneous Localization and Mapping (SLAM) is a core area of study in 3D computer vision, focusing on real-time localization of the camera and generating a fine-grained map of the surrounding environment. It plays a crucial role in robot localization and navigation, autonomous vehicles, and Virtual/Augmented Reality (VR/AR).
Recent breakthroughs in differential rendering, specifically with Neural Radiance Fields (NeRF) [mildenhall2021nerf, martin2021nerf, liu2025i2nerf] and 3D Gaussian Splatting (3DGS) [kerbl20233d, kulhanek2024wildgaussians, liu2024deraings], have significantly advanced visual dense SLAM systems [schops2019bad, huang2021di, runz2017co, bloesch2018codeslam, craig2004tandem, teed2021droid]. NeRF-based neural implicit SLAM [yu2021pixelnerf, barron2021mip, deng2022depth, guo2022nerfren, zhai2024nis] integrates the NeRF model for simultaneous tracking and mapping, which facilitates high-quality, dense online map reconstruction and remarkably improves geometric accuracy. Building on this, Gaussian-based SLAM systems [yan2024gs, keetha2024splatam, matsuki2024gaussian, huang2024photo, peng2024rtg, liu2024structure, deng2024compact] have pushed the boundaries by delivering higher-fidelity map reconstruction. By utilizing explicit Gaussian primitives, 3DGS provides substantial benefits in terms of detailed texture representation [keetha2024splatam], explicit scene manipulation [li2025sgs], and remarkable real-time rendering capabilities [deng2024compact].
@@
However, 3DGS faces particular challenges within the context of SLAM systems. Unlike offline reconstruction often applied to exhaustive image collections [Knapitsch2017, barron2021mip, liu2025realx3d], online SLAM systems are typically deployed in complex environments with insufficient observations, leading to partially observed or obstructed views. This shortage significantly affects the completeness of the Gaussian map, as it struggles to effectively interpolate missing geometry of unobserved areas. Moreover, current Gaussian SLAM systems rely on per-frame pixel backprojection from the input stream, which fails to capture the 3D structure of the scene. Such 2D-based initialization results in a map representation where foreground elements are detailed with denser primitives, while the backgrounds, encompassing broader scene geometries, are less defined. Figure 2 illustrates these limitations, where imbalanced point distribution, erroneous depth projection, and undersampling due to occlusions lead to an uneven and challenging optimization landscape for multiview geometry [niemeyer2024radsplat]. Given these limitations, current Gaussian SLAM systems often maintain a dense keyframe list, such as one in every four frames, during the mapping process that attempts to engage more viewpoints for reliable map optimization [ha2024rgbd]. This reliance on a dense keyframe list demands extensive memory and reduces real-time processing capabilities, both of which are crucial for online SLAM systems.
Figure 2: Gaussian primitives initialized by direct backprojection from RGB-D streams suffer from several drawbacks. The left pixel image illustrates that closer objects occupy more pixels, whereas farther regions receive fewer rays. (a) Imbalanced point sampling occurs due to the uneven distribution of rays. (b) Erroneous projections arise from inaccurate depth estimates in distant regions. (c) Undersampled areas result from occlusions or insufficient views.
To address these challenges, we propose employing NeRF priors to densify the Gaussian SLAM system. This densification is achieved through two primary mechanisms using the NeRF model: (1) its interpolation capabilities efficiently close the gaps in the map with densely positioned Gaussian primitives, and (2) it offers a robust initialization of Gaussian primitives that can be densified using extremely sparse keyframes. Moreover, NeRF-based sampling ensures an even distribution of Gaussian primitives aligned with the scene geometry and allows for manageable granularity in the scene representation through specific sampling ratios. Coupled with loop closure and bundle adjustment (BA), our method delivers state-of-the-art performance across large-scale synthetic and real-world datasets as illustrated in Figure 1. Overall, our contributions can be summarized as follows:
@@ II-A Neural Dense SLAM
In contrast to traditional SLAM systems [mur2015orb, mur2017orb, zhang2019hierarchical, campos2021orb, du2022accurate, chung2023orbeez, pan2024robust] that utilize point-clouds or voxels for sparse map representation, neural dense SLAM systems [yan2017dense, zhang2023go, li2023end, huang2023real, deng2024plgslam, zhu2024nicer, zhou2024mod, deng2024neslam, xie2025depth, liu2025mg] offer substantial advantages through their dense neural radiance maps, providing a robust foundation for downstream tasks in robotics and AR/XR [deng2023prosgnerf]. iMAP [sucar2021imap] pioneered neural implicit SLAM but suffers from large tracking and mapping error using a single MLP. NICE-SLAM [zhu2022nice] employs multiple MLPs for coarser-to-finer mapping, effectively filling the gaps in reconstruction. ESLAM [johari2023eslam] leverages tri-plane features for efficient scene representation, while Co-SLAM [wang2023co] employs multi-resolution hash-grids for real-time performance. Point-SLAM [sandstrom2023point] relies on neural point clouds for dense scene reconstruction, and Loopy-SLAM [liso2024loopy] introduces map corrections to address scene drift caused by accumulated tracking errors. However, maps reconstructed using NeRF typically lack the quality seen in more recent systems built upon 3D Gaussian Splatting. Furthermore, high-resolution NeRF models often require extensive training and exhibit slower real-time rendering performance [kerbl20233d], which significantly reduces their practical efficiency.
@@ II-B Gaussian Splatting SLAM
Propelled by advancements in 3DGS [kerbl20233d], recent Gaussian SLAM systems [yan2024gs, ha2024rgbd, li2025sgs, deng2024compact, hu2025cg] have shown remarkable capabilities in high-fidelity map reconstruction and efficient real-time rendering. Notably, SplaTAM [keetha2024splatam] utilizes isotropic Gaussian representation coupled with dense point-cloud sampling to ensure geometric precision. Conversely, MonoGS [matsuki2024gaussian] employs anisotropic Gaussians to accelerate map reconstruction and enhance texture rendering. Nevertheless, the frame-to-frame tracking mechanisms of these systems do not incorporate loop closure or bundle adjustment, which leads to significant tracking discrepancies in real-world environments. Concurrently, Photo-SLAM [huang2024photo] and RTG-SLAM [peng2024rtg], which integrate feature-based tracking [mur2017orb, campos2021orb] with dense Gaussian maps, achieve superior tracking accuracy and real-time performance. However, they take the trade-off of diminished rendering quality due to the sparse sampling of scene representation. Moreover, systems such as Gaussian-SLAM [yugay2023gaussian] and LoopSplat [zhu2024loopsplat] propose to implement submap division and fusion strategies to tackle the high memory consumption of fine-grained Gaussian maps. While Gaussian SLAM systems offer superior rendering quality compared to NeRF models, the explicit and discrete nature of their scene representation often results in substantial gaps and holes in the reconstructed map due to unobserved or obstructed views commonly seen in online systems. These deficiencies severely affect their efficiency in real-world applications.
@@ II-C Gaussian Splatting with Neural Radiance Prior
Vanilla 3DGS [kerbl20233d] utilizes the sparse point cloud derived from Structure-from-Motion [bregler2000recovering, schonberger2016structure], whose reconstruction quality heavily relies on the accuracy of the initial point cloud [jung2024relaxing]. As an alternative, RadSplat [niemeyer2024radsplat] introduces NeRF model [mildenhall2021nerf] into the Gaussian Splatting framework, aiming to achieve robust real-time rendering of complex real-world scenes. RadSplat [niemeyer2024radsplat] specifically employs NeRF as a prior for initialization and supervision, enabling fast training convergence and enhanced quality.
Compared to RadSplat [niemeyer2024radsplat], our method differs fundamentally in task objectives and methods. RadSplat [niemeyer2024radsplat] focuses on offline reconstruction scenarios where datasets typically consist of 360-degree or dense viewpoint coverage [barron2022mip, Knapitsch2017], making completeness and rendering quality the primary goals. In contrast, our method addresses the challenges of real-time SLAM systems deployed in robotic systems, where the limited sensor field of view and navigation path lead to sparse views and sequential observations. These factors frequently result in unobserved or partially observed geometry due to obstacles, posing significant challenges for explicit Gaussian representations that often leave critical gaps in the reconstruction. To overcome these issues, we leverage the NeRF prior not only for initialization but as a robust interpolation mechanism to interpolate unobserved regions and enable real-time adaptability in environments with sparse views. Moreover, unlike RadSplat [niemeyer2024radsplat] that naively samples one million points from all cast rays, we propose a geometry-aware sampling strategy that operates directly in 3D space. This approach uniformly allocates primitives across object surfaces, effectively mitigating the imbalanced sampling caused by foreground bias.
@@ III-A Neural Radiance Prior
Preliminaries of Neural Radiance Rendering The NeRF model $f_{\theta}$ [mildenhall2021nerf] is a continuous function that predicts colors $\mathbf{c}\in\mathbb{R}^{3}$ and volume density $\sigma\in\mathbb{R}_{+}$ along the sampled rays $\mathbf{r}$. Specifically, given a camera origin $\mathbf{o}\in\mathbb{R}^{3}$ and ray direction $\mathbf{v}$, by uniformly sampling points $\mathbf{x}=\mathbf{o}+\mathbf{d}_{j}\mathbf{v}|_{j\in\{1...\}}$, the pixel color $\mathbf{C}_{\mathrm{N}}$ can be rendered by NeRF using ray marching:
$\mathbf{C}_{\mathrm{N}}=\sum_{j=1}\mathbf{c}_{j}\alpha_{j}T_{Nj},\quad\text{and}\quad T_{Nj}=\prod_{t=1}^{j-1}(1-\alpha_{t})~,$ (1)
@@
NeRF-based Camera Tracking We track the camera-to-world transformation matrix $\mathbf{T}_{wc}\in\mathrm{SE}(3)$ by optimizing the pose using the objective function of NeRF defined in Equation 2. The pose is initialized based on the principle of constant velocity:
$\mathbf{T}_{i}=\mathbf{T}_{i-1}\mathbf{T}^{-1}_{i-2}\mathbf{T}_{i-1}~.$ (3)
To enable tracking, we first bootstrap the NeRF model using a small number of initial frames, whose poses are estimated via the constant velocity assumption. These early frames provide rough supervision to begin learning coarse scene geometry and appearance. Once initialized, the NeRF is incrementally updated as more frames are integrated. For each new frame $i$, a two-stage optimization strategy is adopted following the conventional NeRF-based SLAM system [wang2023co]. We begin with pose-only optimization while keeping the NeRF parameters fixed. After the pose stabilizes, we perform joint optimization of both the camera pose and NeRF weights. The NeRF model is used exclusively for per-frame camera tracking. Once tracking is complete, loop closure and bundle adjustment are applied based on the fine-grained Gaussian map to correct cumulative drift, as detailed in Section III-C.
Geometry-aware Point Sampling Compared to explicit Gaussian representations that require dense viewpoints for thorough scene optimization, NeRF-based models offer remarkable advantages in terms of interpolation, as they can deduce unseen geometry [pumarola2021d, zhu2022nice, yang2023nerfvs]. To harness this capability for efficient real-time performance, we utilize a multi-resolution hash radiance field as proposed by [muller2022instant]. This grid system enables precise interrogation of volumetric data across various resolutions, effectively capturing detailed surface geometries even in sparsely sampled areas. We identify critical surface transitions by setting a density threshold, $\tau_{\rm grid}$, which examines each grid edge to detect transitions where the density at one corner surpasses $\tau_{\rm grid}$, while at its adjacent corner, it does not. At these transitions, we interpolate the surface crossing points using:
@@ III-B Fine-grained Gaussian Map
Multi-scale Gaussian Rendering By initializing through grid sampling from th NeRF model, we represent the scene using a set of anisotropic 3D Gaussian primitives $\{\mathcal{G}_{j}\}$} [kerbl20233d]. Each primitive $\mathcal{G}_{j}$ is defined by a mean $\mu_{j}\in\mathbb{R}^{3}$, a covariance matrix $\Sigma_{j}$, an opacity value $\alpha_{j}\in[0,1]$, coefficients for third-order spherical harmonics $\mathbf{SH}_{j}\in\mathbb{R}^{16}$, and a scaling factor $s_{j}\in\mathbb{R}^{3}$. During the rendering process, these primitives are first projected onto a 2D plane, transforming them into 2D Gaussians. The transformation utilizes a viewing matrix $W$, and the resulting 2D covariance matrix $\Sigma^{\prime}_{j}$ in the image space is computed as follows [zwicker2001surface]:
@@
The radiance field enables interpolated point initialization for Gaussian primitives, yet it can also lead to aliasing effects that diminish the quality of the map during sampling [barron2021mip, barron2023zip]. This issue becomes particularly acute at the edges of objects where small Gaussian floaters result in significant artifacts. Drawing inspiration from [yu2024mip], we adopt a multi-scale Gaussian rendering strategy that consolidates smaller Gaussians into larger ones to enhance scene consistency. Specifically, we use Gaussian functions across four levels of detail, corresponding to down-sampling resolutions of 1×, 4×, 16×, and 64×. Throughout the training phase, we merge smaller, fine-level Gaussians into larger, coarse-level Gaussians. The selection of Gaussians for merging is determined by pixel coverage, which picks Gaussians based on the coverage range set by the inverse of the highest frequency component in the region, denoted as $f_{\rm max}=1/s_{j}$, where $s_{j}$ is the scaling factor.
Ray-Guided Gaussian Pruning Sampling from the NeRF model may also introduce erroneous Gaussian floaters and outliers that affect the quality of the reconstruction. To reduce redundant Gaussians produced by the densification process and enhance rendering efficiency, we implement a pruning strategy based on sampled rays from the NeRF model. Specifically, we use an importance assessment to identify and remove inactive Gaussians from the map during optimization. The importance of each Gaussian is quantified based on its contribution to sampled rays across all input images $\{I_{i}\}_{i=1}^{N}$. Drawing inspiration from [niemeyer2024radsplat], we implement the score function for each primitive as:
$\mathcal{E}(\mathcal{G}_{i})=\max_{r\in I_{i}}(\alpha_{i}^{r}T_{{\mathcal{G}_{i}}}^{r})~,$ (7)
where $\alpha_{i}^{r}T_{\mathcal{G}_{i}}^{r}$ captures the Gaussian $\mathcal{G}_{i}$’s contribution to the final color prediction of a pixel along the ray $r$. We then compute a pruning mask: